OpenStudy (anonymous):

On July 26, 2005, it rained a record 37 inches in Mumbai, India, in a 24-hour period. Which equation of direct variation relates the number of inches of rain y to the number of hours x?

OpenStudy (anonymous):

direct variation means y=kx where k is a constant if we know y is the number of inches and x is the number of hours, we can plug those into y=kx to solve for k the equation will be 37=k*24 k=37/24 the equation will be y=(37/24)x